### v2

@@ 4.2 Quantitative Evaluation
In our comparative analysis with existing virtual try-on methods on VTON HD dataset [4] and DressCode dataset [15], including CP-VTON [19], VITON-HD [4], FS-VTON [9], SDAFN [1], PF-AFN [5], HR-VTON [12], GP-VTON [21], and the current state-of-the-art (SOTA) DCI-VTON [6], our method showcases competitive performance across various metrics (Table 1).In terms of VITON HD [4], our model demonstrates compatibility with this SOTA method while DCI-VTON outperforms previous studies in all evaluated categories. Notably, our approach excels in certain aspects while trailing in others, both in paired and unpaired settings. A significant advantage of our model lies in its remarkable inference speed. As detailed in Table 1, our model achieves the best inference times, which is more than 17.43 times faster than DCI-VTON in 512x384 resolution - 1.01s of ours compared to 17.60s of DCI-VTON. Besides, our method also demonstrates superior performance that is compatible with the DCI-VTON across almost all three subsets of DressCode [15].
Table 1: Quantitative comparison with baselines on VITON HD [4] and DressCode [15]. We multiply KID by 100 for better comparison. * is represented for that we only infer the Tryon model. T(s)↓ is the inference times. Bold and underline denote the best and the second best
Method VITON-HD DressCode - (512 x 384)
@@
DCI-VTON [6] 0.049 0.906 8.02 0.058 0.043 0.896 8.09 0.028 17.60* 0.0301 - 10.82 0.0348 - 12.41 0.0681 - 12.25
Ours 0.056 0.909 7.53 0.07 0.067 0.909 8.43 0.066 1.01 0.0357 0.9495 10.4 0.0417 0.9413 12.69 0.0727 0.886 11.2
@@ 4.3 Ablation Study
Table 2: Condition effectiveness - ablation study on VITONHD (512x384). We multiply KID by 100 for better comparison
Method VITON-HD – 512 x 384
@@
4.3.2 Conditional Mask-aware Post-processing Performance.
@@
Method VITON-HD
@@
4.4 Qualitative Evaluation
In Fig. 6, we showcase composite images from various methods on the VITON-HD Dataset at 512 x 384 resolution. Our approach demonstrates superior performance to DCI-VTON, the current SOTA results on the VITON-HD dataset in terms of detail features. Our architecture generates highly realistic images, surpassing previous studies like VITON-HD and HR-VTON, particularly excelling in costume details. Notably, our model effectively handles challenging cases such as see-through collars - the first example, where DCI-VTON falls short. In intricate scenarios like crossed arms, both DCI-VTON and our model perform well, but our method excels in preserving finer outfit details (the neckline color is white in DCI-VTON and blue in ours). The last row exemplifies DCI-VTON introducing artifacts (the shirt strap in the neck part of the DCI-VTON result), while our approach maintains realism without these artifacts.
Our research not only preserves outfit details and minimizes artifacts but also addresses the crucial issue of retaining identity information. Leveraging our conditional mask-awareness post-processing technique, we successfully address this concern. Figure 7 illustrates the efficacy of our post-processing, comparing results before and after its application alongside the outcomes of the SOTA paper under the same conditions. Our technique effectively preserves identity information, demonstrated in cases where DCI-VTON fails to retain arm tattoos (first and last rows). DCI-VTON introduces artifacts in the neck area (strange necklaces) and the lower-cloth context, compromising image authenticity ($3^{rd}$, $4^{th}$, and $5^{th}$ rows ). Additionally, DCI-VTON struggles to maintain outfit details, as seen in the first line and highlighted regions (purple square) in lines 4 and 5. Our research ensures the maximum retention of identity information for the inference person without compromising output authenticity.
Figure 7: Example for the identity preservation on VTON-HD Dataset (512 x 384). The blue circle represents the important parts that need to be preserved, the green rectangle is for the wrong context or the artifact made by DCI-VTON compared to our approach, and the red rectangle is for the detailed part of the clothes that can not be retained compared to our approach. Please zoom in for better visualization.
5 Conclusions
In this study, we have addressed the challenges of Virtual Try-On technology, by introducing a novel diffusion-based approach that adeptly preserves both garment texture and user identity. Our integrated system, comprising a warping module and a try-on module, enhanced with a mask-awareness post-processing technique, significantly outperforms existing methods in terms of inference speed, being over 17.43 times faster than the current state-of-the-art, while also maintaining superior fidelity in output. This dual focus on efficiency and detail preservation marks a substantial advancement in the field. However, it is important to acknowledge a limitation in our approach: the necessity for an elaborate post-processing process. This additional step, while crucial for ensuring the integrity of the individual’s identity and the garment’s texture, adds complexity to the overall system. Despite this, the proposed method presents a promising solution for real-world applications, offering a more seamless and accurate virtual try-on experience. Future work could aim to streamline this post-processing phase, further enhancing the system’s efficiency and applicability in diverse scenarios, including a wider range of clothing styles and body types.

### v3

@@ 4.2 Quantitative Evaluation
In our comparative analysis with existing virtual try-on methods on VITON HD dataset [4] and DressCode dataset [22], including CP-VTON [28], VITON-HD [4], FS-VTON [9], SDAFN [1], PF-AFN [5], HR-VTON [16], GP-VTON [31], LaDI-VITON [21], StableVITON [14], and the current state-of-the-art (SOTA) DCI-VTON [6], our method showcases competitive performance across various metrics (Table 1). In terms of VITON HD [4], our model demonstrates compatibility with this SOTA method while DCI-VTON outperforms previous studies in all evaluated categories. Notably, our approach excels in certain aspects while trailing in others, both in paired and unpaired settings. A significant advantage of our model lies in its remarkable inference speed. As detailed in Table 1, our model achieves the best inference times, which is more than 17.43 times faster than DCI-VTON in 512x384 resolution, 1.01s of ours compared to 17.60s of DCI-VTON. Besides, our method also demonstrates superior performance that is compatible with the DCI-VTON in almost all three subsets of DressCode [22].
Table 1: Quantitative comparison with baselines on VITON-HD [4] and DressCode [22]. We multiply KID by 100 for better comparison. T(s)↓ is the inference times. Bold and underline denote the best and the second best.
{tblr}
cells = c, cell11 = r=3, cell12 = c=9, cell111 = c=9, cell22 = c=4, cell26 = c=5, cell211 = c=3, cell214 = c=3, cell217 = c=3, vline2,11 = 1, vline2,6,11,14,17 = 2, vline2,6,11,14,17 = 3, vline2,6,11,14,17 = 4-14, hline1,4,11,15 = -, hline2-4 = 2-19, Method & VITON-HD DressCode - (512 x 384)
@@
LaDI-VITON [21] - - - - 0.0986 0.858 12.31 0.567 8.27 0.0654 0.9129 16.18 0.0603 0.9076 16.31 0.1079 0.852 15.80
StableVITON [14] - - - - 0.073 0.888 8.58 0.073 20.58 0.0388 0.937 9.94 - - - - - -
DCI-VTON [6] 0.049 0.906 8.02 0.058 0.043 0.896 8.09 0.028 17.60 0.0301 - 10.82 0.0348 - 12.41 0.0681 - 12.25
FIP-VITON (Ours) 0.056 0.909 7.53 0.07 0.067 0.909 8.43 0.066 1.01 0.0357 0.9495 10.4 0.0417 0.9413 12.69 0.0727 0.886 11.2
@@ 4.3 Ablation Study
Table 2: Condition effectiveness - ablation study on VITON-HD (512x384). We multiply KID by 100 for better comparison.
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, vline2 = 1-7, hline1,3,8 = -, Method & VITON-HD – 512 x 384
@@
4.3.2 Additional Ablation Study.
Since the number of pages is limited, additional ablation experiment parts are shown in the Appendix. We answer the question "Does the advantage come from the modification of the Warping Module or come from the stronger prior?", "What is the trade-off between multiple and single timestep diffusion?", and "What is the impact of modification in the Warping Module?". Furthermore, there is also an ablation experiment to prove the ability of plug-and-play mask-aware postprocessing block.
@@
4.4 Qualitative Evaluation
In Fig. 6, we showcase composite images from various methods on the VITON-HD Dataset at 512 x 384 resolution. Our approach demonstrates superior performance to StableVITON, and DCI-VTON, the current SOTA results on the VITON-HD dataset regarding detail features. Our architecture generates highly realistic images, surpassing previous studies like VITON-HD and HR-VTON, particularly excelling in costume details compared to DCI-VTON and StableVITON. Notably, our model effectively handles challenging cases such as complicated symbols and characters on the clothes the first and third examples, where DCI-VTON and StableVITON fall short. Moreover, in these cases, DCI-VTON introduces artifacts (neck part of the DCI-VTON result), while our approach maintains realism without these artifacts. Furthermore, the second row of Fig. 6 showcases our outperforming in quality in both the scale and reality of the clothes compared with StableVITON and DCI-VTON. The additional qualitative results are shown in the Appendix.
Our research not only preserves outfit details and minimizes artifacts but also addresses the crucial issue of retaining identity information. Leveraging our conditional mask-awareness post-processing technique, we successfully address this concern. Fig. 7 illustrates the efficacy of our post-processing, comparing results before and after its application alongside the outcomes of the SOTA paper under the same conditions. Our technique effectively preserves identity information, demonstrated in cases where DCI-VTON fails to retain arm tattoos (Fig. 1). DCI-VTON introduces artifacts in the neck area (strange necklaces) and the lower-cloth context, compromising image authenticity ($1^{st}$, $2^{nd}$ rows ). Additionally, DCI-VTON and StableVITON struggle to maintain outfit details, $2^{nd}$ and $3^{rd}$ samples in Fig. 7. Our research ensures the maximum retention of identity information for the inference person without compromising output authenticity. The additional qualitative results are shown in the Appendix.
Figure 7: Example for the identity preservation on VTON-HD Dataset (512 x 384). The blue circle represents the important parts that need to be preserved, the green rectangle is for the wrong context compared to our approach, and the red rectangle is for the detailed part of the clothes that can not be retained compared to our approach. Please zoom in for better visualization.
5 Conclusions
In this study, we have addressed the challenges of Virtual Try-On technology, by introducing a novel diffusion-based approach that adeptly preserves garment texture and user identity. Our integrated system, comprising a warping module and a try-on module, enhanced with a mask-awareness post-processing technique, significantly outperforms existing methods in inference speed, being over 17.43 times faster than the current state-of-the-art, while maintaining superior fidelity in output. This dual focus on efficiency and detail preservation marks a substantial advancement in the field. However, it is important to acknowledge a limitation in our approach: the necessity for an elaborate post-processing process. While crucial for ensuring the integrity of the individual’s identity and the garment’s texture, this additional step adds complexity to the overall system. Despite this, the proposed method presents a promising solution for real-world applications, offering a more seamless and accurate virtual try-on experience. Future work could aim to streamline this post-processing phase, further enhancing the system’s efficiency and applicability in diverse scenarios, including a wider range of clothing styles and body types.
Acknowledgments
This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the Grand Information Technology Research Center support program(IITP-2022-2020-0-01489) supervised by the IITP(Institute for Information & communications Technology Planning & Evaluation); and the Institute for Information & Communications Technology Promotion (IITP) grant funded by the Korean government (MSIP) (No. 2022-0-00407).
F Additional Ablation Study
1 Cross-Attention Effectiveness
In this section, we will answer the question "What is the impact of modification in the Warping Module?" by exploring how much the cross-attention mechanism affects the output of the model. Specifically, as mentioned above, we only apply the cross-attention in the stage of feature map resolution smaller than 64x48. In the case of removing the cross-attention block, we replace them with the concatenation operation. In this ablation study, we choose to compare the performance between the warping module with and without the multi-head cross-attention mechanism (W/o Attn) under with and without the conditional post-processing block (W/o Conditional Post-processing). Furthermore, we also measure the effectiveness of the post-processing block for both cases. Based on the result in Table 3, it can be seen that the cross-attention in the warping module provides a better result in almost all the comparison categories. However, applying cross-attention increases the number of parameters by 31% (from 128.4M to 168.5M) with the resolution here being 512x384. We also visualized the qualitative improvement in Figure 8. It is observable that the absence of the cross-attention mechanism can result in distorted outputs from the warping module (as illustrated in the first example of Figure 8), potentially degrading the final visual result. Moreover, in the second example of Figure 8, the collar part of the garment is wrong-warped compared to the cross-attention ones, and the umbrella label also disappears in the non-cross-attention ones. This comparative analysis underscores the critical role of cross-attention in preserving detail and maintaining structural integrity in the warping process, thereby substantiating its inclusion despite the associated increase in computational complexity.
Figure 8: Visual cross-attention ablation studies in our approach. Please zoom in for better quality.
Table 3: Cross attention - ablation study on VITON-HD [4](512x384). We multiply KID by 100 for better comparison. The number of parameters (nParam(M)) is excluded the CLIP vision embedding model
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=5, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, hline1,3,7 = -, Method & VITON-HD – (512 x 384)
LPIPS↓ SSIM↑ FID↓ KID↓ nParam(M)↓
No_Attn + w/o post-processing 0.0691 0.9015 8.55 0.076 128.4
No_Attn + post-processing 0.0670 0.9074 8.52 0.074 128.4
Attn + w/o post-processing 0.0706 0.9016 8.49 0.071 168.5
Attn + post-processing 0.0675 0.9091 8.43 0.066 168.5
2 Noise Level Efficiency - Prior Assessment
This section explores the impact of varying noise levels on the performance of the Try-on Module. To ensure a controlled comparison, we use the same configuration setting for all $\alpha_{n}$ values and the threshold of conditional mask-aware post-processing will still be 0.8 in this ablation study. The noise levels tested in this study are represented by $\alpha_{n}$ values of 2, 5, and 7. As depicted in Table 4, our Try-on Module works best under $\alpha_{n}=5$. In addition, if we add too little noise (low level of noise - $\alpha_{n}=2$), the try-on module performance will decrease massively. Meanwhile, when the level of noise is too high $\alpha_{n}=7$, the Try-on model performance also be affected - slightly reduced. As depicted in Fig. 9, it is easy to see that when the $\alpha_{n}=2$ the skin color output looks not real and several artifacts also appear in this experiment. However, when we increase the value of $\alpha_{n}$ to 5 or 7, the generated image looks more realistic.
Combining with the above assessment, we can answer the question "Does the advantage come from the modification of the Warping Module or come from the stronger prior?" that the proof of our positive impact comes from stronger prior.
Figure 9: Visualize Noise level ablation studies. Please zoom in for better quality.
Table 4: Noise Level ablation study on VITON-HD [4](512x384). We multiply KID by 100 for better comparison
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell82 = c, cell83 = c, cell84 = c, cell85 = c, hline1,3,9 = -, hline2 = 2-5, Method & VITON-HD - (512x384)
LPIPS↓ SSIM↑ FID↓ KID↓
$\alpha_{n}=2$ (w/o post-processing) 0.0924 0.8854 9.54 0.173
$\alpha_{n}=2$ (post-processing) 0.0915 0.8903 9.61 0.175
$\alpha_{n}=5$ (w/o post-processing) 0.0706 0.9016 8.49 0.071
$\alpha_{n}=5$ (post-processing) 0.0675 0.9091 8.43 0.066
$\alpha_{n}=7$ (w/o post-processing) 0.0712 0.9017 8.77 0.095
$\alpha_{n}=7$ (post-processing) 0.0681 0.9091 8.74 0.094
3 Multiple and Single Time-Step Diffusion Trade-Off
Choosing the Single-Time Step as the main approach for the Try-on Module, there is a concern about whether it reduces the quality of the images instead. To answer the question "What is the trade-off between multiple and single timestep diffusion?", we compared our single-step approach and a multiple-step diffusion approach. In this section, we trained our Tryon Module as a vanilla diffusion model with DDIM schedulers 1000 timesteps (DDIM*), and CLIP embedding is added to time embedding, the results of this experiment are shown in Table 5. Our approach has a big gap in all categories. The reason may come from the difference in the objective function. While the DDIM* approach can only apply L2 loss in their noise prediction, ours can use L1, perceptual loss VGG [12], and even adversarial loss (GAN Loss) [13]. Furthermore, the gap in the results also may come from the difference in the time spent for hyper-parameter tuning and the difference in prior.
Table 5: Multiple and Single Time-Step Diffusion Trade-Off - ablation study on VITON-HD [4](512x384). DDIM* is the vanilla diffusion model with DDIM schedulers 1000 timesteps. We multiply KID by 100 for better comparison.
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=5, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, hline1,3,7 = -, Method & VITON-HD – (512 x 384)
LPIPS↓ SSIM↑ FID↓ KID↓ T(s)↓
DDIM* + w/o post-processing 0.247 0.7978 14.05 0.468 460
DDIM* + post-processing 0.241 0.8053 13.93 0.441 460
Ours + w/o post-processing 0.0706 0.9016 8.49 0.071 1.01
Ours + post-processing 0.0675 0.9091 8.43 0.066 1.01
4 Conditional Mask-Aware Threshold
In this analysis, we explore the effectiveness of the threshold value for conditional mask-aware post-processing. It is important to note that the experimental setup remains consistent with the standard implementation, with the only modification being the adjustment of the threshold for the overlapping ratio condition applied during post-processing. Although the ideal theoretical value for this ratio in this algorithm is 1, empirical evidence presented in Table 6 indicates that the most favorable quantitative outcomes are achieved when $R_{overlap}>0.75$, with the second-best performance observed at a threshold ratio of 0.8. However, because the value of 0.75 is too sensitive to the imperfection of the segmentation prediction, we chose the marginally less optimal but more robust threshold of 0.8 for our investigation. Specifically, it is obvious in Fig. 10 that there is an artifact in the arm part (highlighted in red rectangle region) when the threshold value is just 0.75, and there is not if it is 0.8. This observation underscores the delicate balance between achieving optimal overlap ratios and mitigating the risk of artifacts due to segmentation imperfections.
Figure 10: Visualize Conditional Mask-Aware Threshold ablation studies in our approach. Please zoom in for better quality.
Table 6: Overlap ratio $R_{overlap}$ condition for mask-aware post-processing ablation study on VITON-HD [4](512x384). The applying rate AR(%) is represented for the rate of post-processing applied in at least one part of predicted body segmentation in the test set under unpaired configuration. We multiply KID by 100 for better comparison. Bold and underline denote the best and the second best
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=5, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell82 = c, cell83 = c, cell84 = c, cell85 = c, cell86 = c, cell92 = c, cell93 = c, cell94 = c, cell95 = c, cell96 = c, hline1,3,10 = -, hline2 = 2-6, Method & VITON-HD – (512 x 384)
LPIPS↓ SSIM↑ FID↓ KID↓ AR(%)
w/o post-processing 0.0706 0.9016 8.492 0.071 0
$R_{overlap}>0.75$ 0.0673 0.9094 8.423 0.067 83.19
$R_{overlap}>0.80$ 0.0675 0.9091 8.428 0.066 79.63
$R_{overlap}>0.85$ 0.0677 0.9086 8.434 0.067 76.07
$R_{overlap}>0.90$ 0.0680 0.9076 8.452 0.066 70.77
$R_{overlap}>0.95$ 0.0690 0.9053 8.453 0.068 60.14
$R_{overlap}=1$ 0.0706 0.9017 8.464 0.069 17.16
5 Conditional Mask-aware Post-processing Performance.
@@ 4 Conditional Mask-Aware Threshold
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell82 = c, cell83 = c, cell84 = c, cell85 = c, vline2 = -, hline1,3,5,7,9 = -, Method & VITON-HD
@@
G Additional Qualitative Results
1 Results on VITON-HD.
This section presents additional composite images generated by various techniques on the VITON-HD dataset, showcased in Fig. 13 and Fig. 14. The methods compared include VITON-HD [4], HR-VTON [16], DCI-VTON [6], and StableVITON [14]. Our approach surpasses these prior methods, particularly excelling in synthesizing images that faithfully preserve clothing details, notably outperforming DCI-VTON and StableVITON. Evidently, in the first, third, and fourth rows of Figure 14, DCI-VTON inaccurately depicts the context of the clothes collar, a shortcoming effectively addressed by our model. Moreover, instances last row of Figure 13 reveal alterations in clothing style by DCI-VTON, whereas our model consistently maintains output fidelity. DCI-VTON and StableVITON also fail to ensure the detailed features of the garment (colors, symbols, and figures) (as seen in Figure 13 and Figure 14), a challenge effectively addressed by our model.
Additionally, for a comprehensive understanding of the identity preservation aspect, the unmarked version of identities remaining and additional examples are visualized in Figure 18 and Figure 19.
2 Results on DressCode.
In addition to quantitative assessments, we visually showcase the outcomes of our method in Figure 15, Figure 16, and Figure 17. Across these sub-datasets, our method consistently achieves realistic and natural try-on results, effectively preserving unique identity details of the subject, such as muscle structure, tattoos, and necklaces.
Figure 11: Visualization for conditional and unconditional post-processing block. If we use unconditional post-processing for the second one, the final output might get the artifact feature - red rectangle regions.
H Reason for using conditional post-processing.
As mentioned above, we apply the second post-processing block with the condition of overlapping rate. To explain this, we illustrated the result if we used both post-processing blocks unconditionally, the final output can not handle the artifact since the model can not guarantee to generate the same as the real image - as depicted in Figure 11.
I Implemention Details.
For the two main modules of our model, the warping module and the try-on module, we conduct separate training. The warping network is trained for 250 epochs using the Adam optimizer [15] with a learning rate of $5\times 10^{-6}$. Regarding the try-on module, we train it for 500 epochs with the Adam optimizer and a learning rate of $5\times 10^{-5}$. It is noteworthy that for each resolution, we train different configurations for both warping and try-on module networks, and the hardware configuration for training is specified in Table 8, Table 9, and Table 10.
Figure 12: Identity remaining failure.
J Discussion.
1 Conditional Post-processing Limitation.
As mentioned in the ablation study, conditional mask-aware post-processing directly affects the final model performance both qualitatively and quantitatively. Because the overlapping rate is considered the major condition, it greatly relies on the segmentation output of the body part from both the pre-processing human parsing step [17] and our warping module performance. Consequently, the effectiveness of the conditional mask-aware post-processing is inherently tied to the performance of these human parsing predictions. It encounters limitations when dealing with identity information that is situated on the clothing before the virtual try-on process. For example, the identity information lies on the clothing before trying the cloth on, as depicted in Fig. 12. It highlights the need for advanced segmentation and parsing techniques that can more precisely delineate between different types of identity information, ensuring that such details are faithfully preserved and accurately represented in the final output. This area presents an opportunity for further research and development, aiming to enhance the model’s ability to handle a broader range of identity information with greater fidelity.
2 Try-on Module Limitation.
As mentioned, our try-on module only focuses on generating the missing part in the image, fine-tooling the items of clothing, and all the preserved parts remain. Theoretically, this approach only works perfectly under circumstances where the segmentation preprocessing and the warping-based segmentation prediction are executed flawlessly. Meanwhile, that perfection never happens, hence, our diffusion-based module performance is again limited under these segmentation performances. Any inaccuracies or imperfections in the segmentation predictions can lead to errors in the final output, such as misaligned clothing, incorrect texture synthesis, or artifacts in the preserved regions. This limitation underscores the importance of continued advancements in segmentation techniques and the development of more robust models capable of handling imperfections in segmentation predictions.
3 Explicit Warping Module Limitation.
Another limitation of our approach lines on our explicit warping module procedure. This procedure will always generate that appears to have the same fit as the initial clothing the person is wearing. It is because of the use of human parsing as the condition where the preserved mask is used for the input condition.
Figure 13: Qualitative comparison of difference methods on VITON-HD Dataset. Please zoom in for better visualization.
Figure 14: Qualitative comparison of difference methods on VITON-HD Dataset. Please zoom in for better visualization.
Figure 15: Visualization results on DressCode Dataset (512 x 384) - Dress.
Figure 16: Visualization results on DressCode Dataset (512 x 384) - Lower Clothes.
Figure 17: Visualization results on DressCode Dataset (512 x 384) - Upper Clothes.
Figure 18: Examples of unmarked version for the identity preservation on VITON-HD.
Figure 19: Examples of unmarked version for the identity preservation on VITON-HD.
Table 8: Network Configurations
{tblr}
width = colspec = Q[87]Q[263]Q[190]Q[190]Q[204], cells = c, hline1-2,5,12 = -, Module & Parameters VITON-HD (256x192) VITON-HD (512x384) DressCode (512x384)
Warping Number of FPN 5 6 6
Cross-attention resolutions [64,32,16,8] [64,32,16,8] [64,32,16,8]
Cross-attention dropout 0.2 0.2 0.2
Try-on Base channels 256 128 128
Channels multiplier per scale [1, 2, 2, 2, 4] [1, 1, 2, 2, 4] [1, 1, 2, 2, 4]
Attention dropout 0.1 0.1 0.1
Attention resolutions [64,32] [64,32,16] [64,32,16]
Number of ResBlock per scale 2 2 2
CLIP image encoder version ViT-B/32 ViT-B/32 ViT-B/32
Noise level 5 5 5
Table 9: Warping Module Hyper-parameter Configurations
{tblr}
width = colspec = Q[200]Q[204]Q[204]Q[219], columneven = c, column3 = c, cell92 = c=30.627, cell102 = c=30.627, cell112 = c=30.627, cell122 = c=30.627, cell132 = c=30.627, cell142 = c=30.627, hline1-2,9,15 = -, & VITON-HD (256x192) VITON-HD (512x384) DressCode (512x384)
$lr^{w}_{G}$ $5\times e^{-6}$ $5\times e^{-6}$ $5\times e^{-6}$
$lr^{w}_{D}$ $5\times e^{-6}$ $5\times e^{-6}$ $5\times e^{-6}$
Adam (${\beta}_{1}-{\beta}_{2}$) 0.5 - 0.999 0.5 - 0.999 0.5 - 0.999
Batch size 8 4 4
Number of epochs 500 250 250
EMA None None None
Number of GPUs and GPUs 1 RTX 4090 (24GB) 1 A100 Tesla (40GB) 1 A100 Tesla (40GB)
${\alpha}^{w}_{per}$ 0.2
${\alpha}_{ce}$ 3
${\alpha}^{w}_{m}$ 0.3
${\alpha}^{w}_{adv}$ 0.1
${\alpha}_{TV}$ 0.1
${\alpha}_{sec}$ 6
Table 10: Try-on Module Hyper-parameter Configurations
{tblr}
width = colspec = Q[200]Q[204]Q[204]Q[219], columneven = c, column3 = c, cell92 = c=30.627, cell102 = c=30.627, cell112 = c=30.627, hline1-2,9,12 = -, & VITON-HD (256x192) VITON-HD (512x384) DressCode (512x384)
$lr^{tryon}_{G}$ $5\times e^{-5}$ $5\times e^{-5}$ $5\times e^{-5}$
$lr^{tryon}_{D}$ $5\times e^{-5}$ $5\times e^{-5}$ $5\times e^{-5}$
Adam (${\beta}_{1}-{\beta}_{2}$) 0.9 - 0.999 0.9 - 0.999 0.9 - 0.999
Batch size 3 3 3
Number of epochs 500 500 500
EMA 0.9999 0.9999 0.9999
Number of GPUs and GPUs 1 RTX 4090 (24GB) 1 A100 Tesla (40GB) 1 A100 Tesla (40GB)
${\alpha}^{tryon}_{per}$ 1
${\alpha}^{tryon}_{adv}$ 0.1
${\alpha}_{n}$ 5